# InterMask: 3D Human Interaction Generation via Collaborative Masked Modelling

### v1

@@ 5 Conclusion
Limitations and Future Work While InterMask generates high-quality human interactions, there are some limitations. First, when visualized as SMPL Loper et al. (2015) meshes, issues such as body penetration between individuals and jerky movements during rapid actions can occur. Second, our model sometimes interprets motions as dances without explicit prompting, likely due to implicit biases in the training dataset. Lastly, our approach is currently optimized for short sequences up to 10 seconds, reflecting the constraints of our dataset. Visual examples are shown in appendix E. Future research directions include implementing techniques for smoother transitions, addressing dataset biases, and expanding the model’s capacity to handle longer interaction sequences.
Appendix A Implementation Details
@@ A.3 Inference Details
Appendix B Diversity Demonstration
@@
Our quantitative comparison (section 4.1) shows that InterMask prioritizes adherence to text over extreme diversity, while still being able to generate different distinct interactions for the same text prompt. Here, in fig. 9, we show 2 visual examples to demonstrate this. In both cases, the model remains consistent in generating interactions described in the text prompt while exhibiting distinct features in different samples. For the dancing case, first sample shows waving hands in the beginning followed by a synchronized forward step, while individuals face the same direction throughout. The second sample shows individuals facing each other in the beginning, followed by waving hands and a synchronized spin. For the boxing case, first sample shows one individual punching three times and continuously moving forward, while the second sample shows them punching two times and retreating at the end.
Appendix C Alternative Modeling
@@
Appendix D Reaction Generation Inference
@@
Appendix E Failure Cases
@@
Appendix F User Study
We conducted the user study on the Amazon Mechanical Turk platform, where the interface presented to the users is shown in Figure 13. For each sample, users were provided with clear instructions to rate two animations—one generated by InterMask and the other by InterGen (Liang et al., 2024)—from the same text description, with the order of the animations randomized for each sample to avoid bias. Participants were asked to rate both animations on a scale of 1 to 5 for interaction quality and text adherence. Following the individual ratings, they were asked to select the better animation based on their overall impression. A total of 30 samples were evaluated, with each sample being rated by three users. To ensure high-quality feedback, we filtered users to include only those with Amazon Mechanical Turk master status, with a task approval rating of over 97% and more than 1000 previously approved tasks.

### v3

@@ 5 Conclusion
Limitations and Future Work While InterMask generates high-quality human interactions, there are some limitations. First, when visualized as SMPL Loper et al. (2015) meshes, issues such as body penetration between individuals and jerky movements during rapid actions can occur. Second, our model sometimes interprets motions as dances without explicit prompting, likely due to implicit biases in the training dataset. Lastly, our approach is currently optimized for short sequences up to 10 seconds, reflecting the constraints of the dataset. Visual examples are shown in appendix J. Future research directions include implementing techniques for smoother transitions, preventing penetrations, addressing dataset biases, and expanding the model’s capacity to handle longer sequences.
Reproducibility Statement We have made our best effort to ensure reproducibility, including but not limited to: 1) description of our implementation details in appendix E; 2) detailed graphic illustrations of our model architectures, training mechanisms and inference processes in Figures 2, 3, 7, 8, 10 and 13; and 4) public release of code and model checkpoints (appendix A).
Appendix A Supplementary Material
In addition to the subsequent appendix sections, we provide the following as supplementary material for this work:
  • Demo Videos: We provide several demonstration videos on open review, showcasing visual interaction results generated using InterMask. These include an animation gallery with generated interactions for everyday actions, dance and combat; nuanced description demonstration to highlight the capability of Intermask to follow specific details in text; diversity demonstration with multiple generated interactions for each text prompt; comparison videos to compare the generated results of InterMask and InterGen (Liang et al., 2024); comparison videos for the ablation study on Inter-M Transformer; an animation gallery to show reaction generation results; some results on more complex prompts; some results showing longer 10 second interactions; and finally some failure cases and some results addressing the concerns on fluidity of the generated motions. The videos are shown in form of a webpage, provided as an html file.
  • Code Implementation: We provide the open-source code implementation of our method to ensure reproducibility.
[…]
    github.com/gohar-malik/intermask
Appendix B Motion VQ-VAE
Our VQ-VAE framework, illustrated in Figure 7, constructs a 2D motion token map to represent individual motion sequences in a discrete manner, while retaning both spatial and temporal dimensions. The encoder processes motion sequences represented as $\mathbf{m}_{p}\in\mathbb{R}^{N\times J\times d}$, where $N$ is the number of poses, $J$ is the number of joints, and $d$ is the joint feature dimension. By employing 2D convolutional layers, the encoder effectively captures spatial and temporal dependencies within the motion data while progressively downsampling both dimensions. The downsampling process is achieved through strided 2D convolutions and ResNet blocks, which also use 2D convolutions along with dropout layers. This results in a latent representation of size $\mathbf{\tilde{t}_{p}}\in\mathbb{R}^{n\times j\times d^{\prime}}$, where $n$ and $j$ are the downsampled temporal and spatial dimensions, and $d^{\prime}$ is the latent feature dimension.
The latent representation $\mathbf{\tilde{t}_{p}}$ is quantized using a learned codebook $\mathcal{C}$ with $|\mathcal{C}|$ entries. Each feature vector $\tilde{t}_{i}$ in $\mathbf{\tilde{t}_{p}}$ is replaced by the index of its nearest codebook entry, using the vector quantization process:
$\mathbf{q}(\tilde{t}_{i})=\arg\min_{c_{k}\in\mathcal{C}}\|\tilde{t}_{i}-c_{k}\|^{2},$ (9)
where $c_{k}$ represents the codebook entries.
The resulting quantized representation is a 2D motion token map $t_{p}$, where each token encodes local spatio-temporal context. This design enables the preservation of both spatial and temporal dimensions in the motion data, enhancing the model’s ability to generate realistic and contextually accurate interactions.
Figure 7: Detailed illustration of the 2d discrete motion token map construction. The 2d encoder, consisting of 2d convolutional layers, downsamples the input motion from $(N,J,d)$ to $(n,j,d^{\prime})$. The downsampled representation is then quantized by replacing each vector with the index of its closest vector in the learned codebook.
Appendix C VQ-VAE Geometric Losses
Equation 10 shows the geometric losses used to train our Motion VQ-VAE to impose physical and geometric constraints on the reconstructed motion, in a data-driven way. The velocity loss $\mathcal{L}_{vel}$ encourages the reconstructed motion sequences to obey the velocity of joints in the ground truth sequences and the bone length loss $\mathcal{L}_{bl}$ the distance between adjacent joints. The foot contact loss $\mathcal{L}_{f}c$ encourages the feet to have zero velocity whenever they are in contact with the ground.
Appendix D Two-Stage Token Masking
Interaction Generation Reaction Generation
$p_{r}$ FID $\downarrow$ R Prec (Top1) $\uparrow$ FID $\downarrow$ R Prec (Top1) $\uparrow$
0.7 5.214 0.447 2.850 0.476
0.8 5.154 0.449 2.991 0.462
0.9 5.152 0.450 3.368 0.416
Table 4: Interaction Generation and Reaction Generation results of different values of $p_{r}$. Bold face indicates the best result, while underscore refers to the second best.
Figure 8: Illustration of the two-stage masking technique used during training of the Inter-M Transformer. For stage 1, we either apply Random Masking with a probability of $p_{r}$ or Interaction Masking with probability $1-p_{r}$. In stage 2, we apply the Step Unroll Masking on the predicted tokens from stage 1.
Here we provide more details on the masking strategy (section 3.2.1) used in training the Inter-M Transformer. As illustrated in Figure 8, the two-stage masking technique begins with random masking or interaction masking in the first stage. Random masking teaches the model to predict random tokens from both individuals. Whereas Interaction masking, where only one individual’s tokens are masked, promotes learning inter-person dependencies critical for coherent interactions and to improve performance in the reaction generation task. The masking strategy alternates between these two methods based on a probability parameter $p_{r}$, with random masking applied $p_{r}$ of the time and interaction masking $(1-p_{r})$. To evaluate the effect of $p_{r}$, we test different values ${0.7,0.8,0.9}$ and find that $p_{r}=0.8$ offers the best balance between interaction and reaction generation, as shown in Table 4. In the second stage, step unroll masking is applied which retains some of the predicted tokens from stage 1, remasks the remaining tokens and predicts them again. This is employed to incorporate the inference-time progressive refinement of tokens in the training process.
Appendix E Implementation Details
@@ E.2 Training Details
$p_{r}$ 0.8 Random Masking probability for stage 1 masking during training
@@ E.3 Inference Details
Appendix F Diversity Demonstration
@@
Our quantitative comparison (section 4.1) shows that InterMask prioritizes adherence to text over extreme diversity, while still being able to generate different distinct interactions for the same text prompt. Here, in Figure 9, we show 2 visual examples to demonstrate this. In both cases, the model remains consistent in generating interactions described in the text prompt while exhibiting distinct features in different samples. For the dancing case, first sample shows waving hands in the beginning followed by a synchronized forward step, while individuals face the same direction throughout. The second sample shows individuals facing each other in the beginning, followed by waving hands and a synchronized spin. For the boxing case, first sample shows one individual punching three times and continuously moving forward, while the second sample shows them punching two times and retreating at the end.
Appendix G Alternative Modeling
@@
Appendix H Ablation Study Qualitative Results
Figure 12: Qualitative results for the ablation study on Inter-M Transformer to verify contributions of the proposed Attention modules.
In this section, we present qualitative results for our ablation studies to complement the quantitative findings discussed in section 4.2. Figure 11 shows two ground truth interaction sequences with their reconstructed samples from the 2D token map VQ-VAE and the baseline 1D token map VQ-VAE. As shown, the 1D VQ-VAE struggles to accurately reconstruct the spatial positions and orientations of the joints for both individuals, leading to incorrect positioning and orientation relative to each other. This results in not only unrealistic interactions but also bizarre individual poses. In contrast, the proposed 2D VQ-VAE provides highly accurate reconstructions at both the individual pose and the collective interaction level.
Figure 12 illustrates the impact of different attention modules in our Inter-M transformer by comparing outputs by removing the spatio-temporal attention, cross-attention, and self-attention modules independently. We provide results for three distinct interaction scenarios: boxing, sneaking, and synchronized dancing. The spatio-temporal attention module emerges as a critical component for generating complex poses and ensuring spatial awareness in interactions. Without this module, the boxing scenario exhibits overly simplistic poses, such as timid punching and blocking, alongwith the individuals failing to face each other properly. In the sneaking scenario, the absence of spatio-temporal attention eliminates the essential spatial progression, as the sneaking individual does not gradually approach the other. Similarly, in the dancing scenario, the generated motions are reduced to basic hand raises, and one individual adopts an unnatural crouching pose. By contrast, the inclusion of spatio-temporal attention enables accurate spatial positioning and expressive, synchronized interactions. The cross-attention module seems vital for modeling inter-person dependencies, particularly in achieving accurate reaction timing. Without it, the response motions of the interacting individual appear either delayed or prematurely executed across all examples. For instance, in the boxing scenario, the reactive movements fail to synchronize with the initiating individual’s punches. In the dancing scenario, the lack of cross-attention results in poor synchronization, disrupting the fluidity of the interaction. Lastly, the self-attention module serves as a refinement mechanism, enhancing the overall quality and coherence of individual motions. Its removal introduces subtle inconsistencies, such as jerky transitions or less fluid movements, which slightly degrade the interaction’s realism.These observations collectively underscore the importance of each attention module in generating realistic, contextually accurate, and expressive interactions.
Appendix I Reaction Generation Inference
@@
Appendix J Failure Cases
@@
Appendix K User Study
We conducted the user study on the Amazon Mechanical Turk platform, where the interface presented to the users is shown in Figure 15. For each sample, users were provided with clear instructions to rate two animations—one generated by InterMask and the other by InterGen (Liang et al., 2024)—from the same text description, with the order of the animations randomized for each sample to avoid bias. Participants were asked to rate both animations on a scale of 1 to 5 for interaction quality and text adherence. Following the individual ratings, they were asked to select the better animation based on their overall impression. A total of 16 users evaluated a total 30 samples, with each sample being rated by three users. To ensure high-quality feedback, we filtered users to include only those with Amazon Mechanical Turk master status, with a task approval rating of over 97% and more than 1000 previously approved tasks.